The nuclei of which one of the following pairs of nuclei are isotones?


(a)    34Se74, 31Ga71


(b)    38Sr84, 38Sr86


(c)    42Mo92, 40Zr92


(d)    20Ca40, 16S32


Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

(a)

Isotones means equal number of neutrons i.e.,

(A-Z) = 74 - 34 = 71 - 31 = 40.